# GRAM: Spatial general-purpose audio representation models for real-world applications

### v3

@@ Abstract
Although audio foundations models have seen great progress on a wide variety of tasks, their application in real-world acoustic environments with reverberation and noise has been less successful. Moreover, as audio foundation models are typically trained on dry, single-channel audio clips, the inherent spatial nature of real-world sound scenes is overlooked and tasks involving sound localization ruled out. To address these limitations, we propose GRAM: a General-purpose Real-world Audio Model utilizing a multi-channel masked auto-encoder approach to efficiently learn spatial audio representations from high-quality simulated real-world scenes. To evaluate the performance of GRAM and other audio foundation models in real-world sound scenes, we release Nat-HEAR: A naturalistic version of the HEAR benchmark suite comprising a simulated real-world version, as well as two new sound localization tasks. We show that the performance of GRAM surpasses all state-of-the-art self-supervised audio foundation models and speech models on both HEAR and Nat-HEAR, while using only a fraction of the training data. GRAM also showcases state-of-the-art localization performance, surpassing even supervised sound localization approaches, and can be flexibly applied either to a two-channel, binaural sound format or a four-channel, Ambisonics format. Validating GRAM’s performance on real-world sound recordings demonstrates robust transfer to real-world scenes. Taken together, GRAM presents a significant advancement towards robust, spatial audio foundation models for real-world applications.
@@ 1 Introduction
Despite the complexity and diversity of everyday sound scenes, human listeners effortlessly interact with their acoustic environment in myriad ways. Audio foundation models that perform a similar, human-like range of tasks have received widespread attention (Turian et al., 2022; Wang et al., 2022a; Yang et al., 2021). While these models demonstrate strong performance on audio benchmarks with minimal fine-tuning (for example, (Chen et al., 2023; Baevski et al., 2020; Yadav et al., 2024), they overlook inherent aspects of real-world sound scenes: the spatial dimension, reverberation and background noise. Specifically, audio foundation models are typically trained on large-scale sound datasets consisting of dry, non-spatial sound clips such as AudioSet (Gemmeke et al., 2017) and Librispeech (Panayotov et al., 2015). The effectiveness of these approaches in naturalistic, complex acoustic environments with background noise and reverberation is therefore limited.
@@
Crucially, the lack of spatial information in audio embeddings precludes sound localization tasks and the use of spatial sound features for improving performance on complex audio tasks such as audio scene analysis. Audio scene analysis refers to the separation of overlapping sound waves in complex multi-source sound scenes and the subsequent grouping of the frequency components into coherent auditory objects (Bregman, 1984; Bizley & Cohen, 2013). In humans, such audio scene analysis is aided by spatial cues as well (Bizley & Cohen, 2013; van der Heijden et al., 2019). Similarly, incorporating spatial knowledge into universal audio embedding models is expected to benefit downstream tasks where ambient intelligence and acoustic awareness are desired, such as acoustic scene understanding.
@@
A major challenge for the development of audio foundation models for real-world sound scenes is the scarcity of naturalistic sound data for model training. Recording a vast amount of spatial sound scenes with varying reverberation characteristics is infeasible and requires fine-tuned recording setups and conditions (Zheng et al., 2024). The simulation of spatial acoustic scenes has therefore received much attention and progressed from simulating shoebox rooms (Scheibler et al., 2018) to simulating realistic sound scenes from everyday life (Chen et al., 2020). Yet, no large-scale datasets of naturalistic sound scenes exist to date, hampering both the development as well as the systematic evaluation of audio foundation models for real-world sound scenes. For example, benchmark task suites such as HEAR (Turian et al., 2022), HARES (Wang et al., 2022a) and SUPERB (Yang et al., 2021) solely contain datasets consisting of dry, non-spatial sound clips without background noise and do not include sound localization tasks.
@@
To address these limitations of audio foundation models for real-world applications, we present GRAM (General-purpose, Real-world Audio Model). GRAM is a self-supervised, multi-channel masked auto-encoder model that efficiently learns spatial general-purpose audio representations from simulated real-world sound scenes. To train GRAM, we developed a custom pipeline which makes use of the Soundspace 2.0 platform (Chen et al., 2022a) to simulate high-quality real-world sound scenes from AudioSet (Gemmeke et al., 2017), and of WHAMR! (Maciejewski et al., 2020) for adding background noise. Further, to promote the systematic evaluation of audio foundation models on naturalistic sound scenes, we introduce Nat-HEAR. Nat-HEAR is an extension of the HEAR benchmark suite which contains simulated real-world versions of the downstream tasks, and additionally includes two sound localization tasks. We experiment with two state-of-the-art encoder architectures (Transformer and Mamba) to assess which architecture is most suitable for spatial general-purpose audio representation learning in our multi-channel masked auto-encoder approach. We present two versions of GRAM to ensure flexible application across audio formats: GRAM-Binaural for two-channel audio clips, and GRAM-Ambisonics for four-channel audio clips in the first-order Ambisonics format. Finally, we perform systematic ablation experiments on mask type (patch versus time-based), ratio of simulated real-world sound scenes and conventional dry sound clips in pre-training, mask ratio, and in-batch sampling to elucidate which factors are critical for successful spatial general-purpose audio representation learning. [^1] All code and materials are available at http://tbd.
@@
Nat-HEAR: To encourage systematic evaluation of audio foundation models on naturalistic scenes, we present an extended version of the HEAR benchmark suite (Turian et al., 2022) in which we transform the HEAR datasets in the HEAR downstream tasks to naturalistic versions. Additionally, we add two novel, naturalistic sound localization tasks in Nat-HEAR.
@@ 2 Related Work
Supervised audio representation learning: Supervised methods for audio representation learning have achieved notable success in recent years. Approaches such as AST (Gong et al., 2021a), PaSST (Koutini et al., 2022) and HTS-AT (Chen et al., 2022b) have Transformer-based architectures as a backbone, for example ViT (Dosovitskiy et al., 2021) and Swin Transformer (Liu et al., 2021). To mitigate the need for large annotated datasets, some of these approaches are based on models pre-trained on image data (for example, AST (Gong et al., 2021a), PSLA (Gong et al., 2021b)). Question and answer models constitute a more recent category of supervised approaches that integrate audio representation learning with large language models (for example, Spatial-AST (Zheng et al., 2024) and Qwen-Audio (Chu et al., 2023). However, supervised training requires large-scale annotated datasets and is sub-optimal for learning general-purpose audio representations that generalize across tasks.
Self-supervised audio representation learning: Self-supervised audio representation learning approaches aim to learn robust audio representations that generalize to a wide variety of tasks (Wang et al., 2022a; Turian et al., 2022). Masking-based approaches utilizing transformer backbones to reconstruct masked patches of input spectrograms currently constitute pre-dominant approach, including (SSAST (Gong et al., 2022)), MSM-MAE (Niizumi et al., 2022), MaskSpec (Chong et al., 2023), MAE-AST (Baade et al., 2022) and Audio-MAE (Huang et al., 2022). Of the masked auto-encoder approaches, MW-MAE (Yadav et al., 2024) achieves state-of-the-art performance on the HEAR benchmark by using multi-window local-global attention in the decoder. Recently, SSAM (Yadav & Tan, 2024) utilized a Mamba (Gu & Dao, 2023) architecture in their encoder and achieved similar performance as MW-MAE. In contrast to the masked auto-encoders, BEATS (Chen et al., 2023) utilizesmasking-based approach based on latent embeddings extracted by an acoustic tokenizer. Finally, successful self-supervised approaches that do not rely on masking at all include contrastive learning frameworks such as COLA (Saeed et al., 2021).
Another category of self-supervised audio representation models focuses on speech representations specifically, making use of generative, predictive or contrastive learning (Mohamed et al., 2022). These speech models are typically trained on datasets such as Librispeech (Panayotov et al., 2015) or LibriLight (Kahn et al., 2020) and include state-of-the-art models such as Wav2Vec2 (Baevski et al., 2020), HuBERT (Hsu et al., 2021) and WavLM (Chen et al., 2021). However, while these models excel at speech-based tasks, they do not necessarily generalize well to non-speech sounds and non-speech tasks (Turian et al., 2022). Crucially, none of the existing self-supervised approaches for audio or speech representation learning optimize for performance in real-world sound scenes that are spatial, reverberant, and noisy.
@@ 3.1 Simulating real-world acoustic scenes
Pipeline overview: A room impulse response (RIR) captures room specific acoustic properties such as reverberation. We utilized high-resolution, detailed 3D meshes of houses with various architectural characteristics from Matterport3D (Chang et al., 2017) in order to simulate RIRs for many different rooms in each house with the Monte Carlo ray tracing RIR simulator provided by SoundSpaces 2.0 Chen et al. (2022a). SoundSpaces 2.0 combines the simulated RIRs with a head-related transfer function (HRTF) (Algazi et al., 2001) to generate a binaural RIR (BRIR) or with an ambisonics microphone configuration to generate an ambisonics RIR (ARIR). BRIRs capture both room acoustic properties and human spatial hearing characteristics introduced by the shape of the ears, head and torso, while ARIRs capture room acoustic properties as well as the spatial cues encoded in first-order Ambisonics.
@@ 3.2 GRAM framework
The GRAM learns spatial audio representation by reconstructing multi-channel masked spectrogram patches. First, a patch extractor consisting of a single convolutional layer with 2D convolutional filters divides each multi-channel spectrogram into $n$ non-overlapping patches $P_{1},\ldots,P_{n}$ with $P_{i}\in\mathbb{R}^{C\times T\times F}$, and embeds each patch into a linear patch embedding $E_{i}\in\mathbb{R}^{768}$ (Figure 1). Non-masked patch embeddings are input to the encoder, for which we selected the 12-layer ViT-Base (ViT-B) Transformer (Dosovitskiy et al., 2021) similar to Huang et al. (2022); Yadav et al. (2024). The encoder outputs patch representations $O_{i}\in\mathbb{R}^{768}$ for $i=1,\ldots,n$, where $n$ is the number of unmasked patches. Finally, a Transformer decoder with local-global attention (Yadav et al., 2024) followed by a linear head takes all patch representations $O_{1},\ldots,O_{n}$ as well as all masked patches $M_{1},\ldots,M_{n}$ to reconstruct the multi-channel spectrogram from last layer embeddings.
@@ 3.3 Pre-training
Online mixing of naturalistic sound scenes: The 85,000 naturalistic scenes were split into a train set of 70,000 scenes (corresponding to 70 Matterport3D houses), and a test set of 15,000 scenes (15 Matterport3D houses) for down-stream evaluation (see Section 3.4). We used the 70,000 naturalistic scenes in the train set to generate naturalistic scenes for all audio clips in the unbalanced training set of AudioSet (10-second sound tracks of 1.74 million YouTube videos (Gemmeke et al., 2017)). Specifically, during training we randomly paired an AudioSet clip with a noise sound clip from the WHAMR! background noise database (Maciejewski et al., 2020). WHAMR! noise clips longer than 10 s were trimmed to 10 s duration and a linear fade-in/fade-out of 200 ms was applied to every noise clip prior to mixing of the sound scene.
@@
Patch extraction and masking: For pre-training, we divided the binaural spectrogram into $P_{i}\in\mathbb{R}^{2\times 8\times 16}$, and ambisonics spectrograms into $P_{i}\in\mathbb{R}^{7\times 8\times 16}$ patches. We used an adapted version of the mask-based framework of MW-MAE (Yadav et al., 2024), randomly selecting a subset of $n$ patches $M_{1},\ldots,M_{n}$ for $i=1,\ldots,n$ for masking (masking ratio = 0.8) and replacing their embedding with a learnable mask token. Finally, we added fixed sinusoidal positional embeddings to all embedded patches.
Decoder with local-global attention: The decoder takes as input both the unmasked patches $O_{1},\ldots,O_{n}$ with $O_{i}\in\mathbb{R}^{768}$, and the masked patches $M_{1},\ldots,M_{n}$ with $M_{i}\in\mathbb{R}^{768}$ as well as fixed sinusoidal positional embeddings for each patch (Figure 1). To implement local-global attention (Yadav et al., 2024), we selected window sizes of [2, 5, 10, 25, 50, 100, 0, 0]. Here, 0 signifies global plain attention.
Pre-training specification: We trained all GRAMs for 500 K steps on an H100 92 GB GPU machine with 16 CPU cores. We used the AdamW optimizer (Loshchilov & Hutter, 2017) with weight decay rate of 0.01, gradient clipping, and a cosine learning rate scheduler with 10 K steps warm-up. The initial learning rate was set to 0.0002, and decayed to 0. We optimize the mean squared error (MSE) loss function between the predicted masked patches and their corresponding input spectrogram patches.
@@ 3.4 Experiments
Model comparison: We compare the performance and efficiency of GRAM-Binaural, GRAM-Ambisonics and GRAM-Clean (trained on regular AudioSet) on downstream tasks with state-of-the-art self-supervised audio representation models with a similar number of parameters as GRAM (90 M): MAE-16x16 (Huang et al., 2022), SSAST-patch (Gong et al., 2022), BEATs-iter3 (Chen et al., 2023), MW-MAE-B-200-4x16 (Yadav et al., 2024), SSAM (Yadav & Tan, 2024); self-supervised speech representation models Wav2Vec 2.0 Base (Baevski et al., 2020), HuBERT Base (Hsu et al., 2021), WavLM Base (Chen et al., 2021).
@@
Downstream tasks: We evaluate GRAM and other state-of-the-art models on the HEAR benchmark task suite, which presents a wide range of tasks to evaluate the downstream performance of audio representation models (Turian et al., 2022). To avoid redundancy we selected the same subset of HEAR tasks as previously used in (Yadav et al., 2024). To enable in-depth evaluation of audio scene analysis capabilities, we added the time-stamp based sound event detection task DCASE-2016 Task 2 (Mesaros et al., 2018) from the HEAR benchmark suite.
@@
We additionally evaluated performance on simulated real-worldsound scenes using Nat-HEAR, which provides a naturalistic version of all selected datasets in the HEAR benchmark suite in two audio formats: a two-channel, binaural format and a four-channel, first-order Ambisonics format. Weincluded sound localization tasks for two different domains which we generated using HEAR benchmark datasets: A speech localization task based on SC-5, and an environmental sound localization task based on ESC-50. The localization tasks are modeled as a multi-output regression task in which model outputs represent the estimated 3D Cartesian coordinates $[x,y,z]$ on the unit sphere (Adavanne et al., 2018). Finally, to assess the transferability of GRAM to real-world sound scenes, we evaluate also on the sound event detection and localization tasks in TUT Sound Events 2018 REAL (Adavanne et al., 2019).
@@
Downstream evaluation: To evaluate the single-channel SOTA audio representation models on NatHEAR, we utilized the omnidirectional channel $W$ of the first-order Ambisonics (Zotter & Frank, 2019) version of Nat-HEAR as model input. The outputs of MAE, SSAST and BEATs were not suitable for the time-stamp based DCASE-2016 sound event detection task. Hence, these models were not evaluated on the DCASE-2016 task. Further, we included Spatial-AST (Zheng et al., 2024) as this is the only model trained on spatial sound scenes and therefore the sole model evaluated on the Nat-HEAR localization tasks besides GRAM-Binaural and GRAM-Ambisonics. To evaluate two-channel models such as GRAM-Binaural on HEAR, we duplicated the single-channel spectrograms of the original HEAR to generate model compatible input. Following the HEAR protocol (Turian et al., 2022) for downstream task evaluation, we extracted embeddings from the frozen pretrained models and trained a shallow downstream classifier on these embeddings to assess how well the learned representations generalize to a broad range of tasks. The embedding extraction for GRAM is described in Appendix D.
@@
Figure 2: Downstream model performance on naturalistic sound scenes. (A) Boxplots of the difference in performance on HEAR and Nat-HEAR, excluding the DCASE-2016 task. Box limits reflect the first and third quartile, center line the median. (B) Circles reflect the difference in performance on HEAR and Nat-HEAR per task. Circle size reflects performance on HEAR (see also Table 5). (C) Nat-HEAR and HEAR downstream performance as a function of training data quantity, (D) Polar plot depicting the performance drop of the three implementations of GRAM per Nat-HEAR task.
@@
Quantifying overall performance: We calculate for each model $m$, the score $s(m)$ to give an impression of the overall performance, similar to (Yang et al., 2021). This score reflects a model’s improvement with respect to the maximum improvement over the baseline obtained by the current state-of-the-art model, averaged across all tasks included in the benchmark. This metric effectively ranks the improvement of models over the baseline as a function of the current maximum improvement (see Appendix D). We use the HEAR-Naive baseline based on mel-spectrograms (Turian et al., 2022).
@@
Generalization to recorded real-world scenes: To evaluate GRAM-Ambisonics on the sound event detection and sound localization takss of TUT Sound Events 2018 REAL (Adavanne et al., 2019), resampled waveforms to 32 kHz and extracted segments with their respective localization labels. We modeled the localization task as polar coordinate regression, predicting [azimuth, elevation], and used the same fine-tuning protocol as for HEAR tasks.

### v4

@@ Abstract
Although audio foundations models have seen great progress on a wide variety of tasks, their application in real-world acoustic environments with reverberation and noise has been less successful. Moreover, as audio foundation models are typically trained on dry, single-channel audio clips, the inherent spatial nature of real-world sound scenes is overlooked and tasks involving sound localization ruled out. To address these limitations, we propose GRAM: a General-purpose Real-world Audio Model utilizing a multi-channel masked auto-encoder approach to efficiently learn spatial audio representations from high-quality simulated real-world scenes. To evaluate the performance of GRAM and other audio foundation models in real-world sound scenes, we release Nat-HEAR: A naturalistic version of the HEAR benchmark suite comprising a simulated real-world version, as well as two new sound localization tasks. We show that the performance of GRAM surpasses all state-of-the-art self-supervised audio foundation models and speech models on both HEAR and Nat-HEAR, while using only a fraction of the training data. GRAM also showcases state-of-the-art localization performance, surpassing even supervised sound localization approaches, and can be flexibly applied either to a two-channel, binaural sound format or a four-channel, Ambisonics format. Validating GRAM’s performance on real-world sound recordings demonstrates robust transfer to real-world scenes. Taken together, GRAM presents a significant advancement towards robust, spatial audio foundation models for real-world applications.  [^1] All code and materials are available on https://github.com/labhamlet, and https://huggingface.co/labhamlet
@@ 1 Introduction
Despite the complexity and diversity of everyday sound scenes, human listeners effortlessly interact with their acoustic environment in myriad ways. Audio foundation models that perform a similar, human-like range of tasks have received widespread attention (HEAR; HARES; superb). While these models demonstrate strong performance on audio benchmarks with minimal fine-tuning (for example, (beats; wav2vec2; mwmae), they overlook inherent aspects of real-world sound scenes: the spatial dimension, reverberation and background noise. Specifically, audio foundation models are typically trained on large-scale sound datasets consisting of dry, non-spatial sound clips such as AudioSet (audioset) and Librispeech (librispeech). The effectiveness of these approaches in naturalistic, complex acoustic environments with background noise and reverberation is therefore limited.
@@
Crucially, the lack of spatial information in audio embeddings precludes sound localization tasks and the use of spatial sound features for improving performance on complex audio tasks such as audio scene analysis. Audio scene analysis refers to the separation of overlapping sound waves in complex multi-source sound scenes and the subsequent grouping of the frequency components into coherent auditory objects (bregman1984auditory; bizley2013and). In humans, such audio scene analysis is aided by spatial cues as well (bizley2013and; heijden2019cortical). Similarly, incorporating spatial knowledge into universal audio embedding models is expected to benefit downstream tasks where ambient intelligence and acoustic awareness are desired, such as acoustic scene understanding.
@@
A major challenge for the development of audio foundation models for real-world sound scenes is the scarcity of naturalistic sound data for model training. Recording a vast amount of spatial sound scenes with varying reverberation characteristics is infeasible and requires fine-tuned recording setups and conditions (spatialast). The simulation of spatial acoustic scenes has therefore received much attention and progressed from simulating shoebox rooms (pyroomacoustics) to simulating realistic sound scenes from everyday life (chen20soundspaces). Yet, no large-scale datasets of naturalistic sound scenes exist to date, hampering both the development as well as the systematic evaluation of audio foundation models for real-world sound scenes. For example, benchmark task suites such as HEAR (HEAR), HARES (HARES) and SUPERB (superb) solely contain datasets consisting of dry, non-spatial sound clips without background noise and do not include sound localization tasks.
@@
To address these limitations of audio foundation models for real-world applications, we present GRAM (General-purpose, Real-world Audio Model). GRAM is a self-supervised, multi-channel masked auto-encoder model that efficiently learns spatial general-purpose audio representations from simulated real-world sound scenes. To train GRAM, we developed a custom pipeline which makes use of the Soundspace 2.0 platform (chen22soundspaces2) to simulate high-quality real-world sound scenes from AudioSet (audioset), and of WHAMR! (Maciejewski2020WHAMR) for adding background noise. Further, to promote the systematic evaluation of audio foundation models on naturalistic sound scenes, we introduce Nat-HEAR. Nat-HEAR is an extension of the HEAR benchmark suite which contains simulated real-world versions of the downstream tasks, and additionally includes two sound localization tasks. We experiment with two state-of-the-art encoder architectures (Transformer and Mamba) to assess which architecture is most suitable for spatial general-purpose audio representation learning in our multi-channel masked auto-encoder approach. We present two versions of GRAM to ensure flexible application across audio formats: GRAM-Binaural for two-channel audio clips, and GRAM-Ambisonics for four-channel audio clips in the first-order Ambisonics format. Finally, we perform systematic ablation experiments on mask type (patch versus time-based), ratio of simulated real-world sound scenes and conventional dry sound clips in pre-training, mask ratio, and in-batch sampling to elucidate which factors are critical for successful spatial general-purpose audio representation learning.
@@
Nat-HEAR: To encourage systematic evaluation of audio foundation models on naturalistic scenes, we present an extended version of the HEAR benchmark suite (HEAR) in which we transform the HEAR datasets in the HEAR downstream tasks to naturalistic versions. Additionally, we add two novel, naturalistic sound localization tasks in Nat-HEAR.
@@ 2 Related Work
Supervised audio representation learning: Supervised methods for audio representation learning have achieved notable success in recent years. Approaches such as AST (ast), PaSST (koutini22passt) and HTS-AT (hts) have Transformer-based architectures as a backbone, for example ViT (dosovitskiy2020vit) and Swin Transformer (liu2021Swin). To mitigate the need for large annotated datasets, some of these approaches are based on models pre-trained on image data (for example, AST (ast), PSLA (psla)). Question and answer models constitute a more recent category of supervised approaches that integrate audio representation learning with large language models (for example, Spatial-AST (spatialast) and Qwen-Audio (chu2023qwen). However, supervised training requires large-scale annotated datasets and is sub-optimal for learning general-purpose audio representations that generalize across tasks.
Self-supervised audio representation learning: Self-supervised audio representation learning approaches aim to learn robust audio representations that generalize to a wide variety of tasks (HARES; HEAR). Masking-based approaches utilizing transformer backbones to reconstruct masked patches of input spectrograms currently constitute pre-dominant approach, including (SSAST (gong2022ssast)), MSM-MAE (masked-modelling), MaskSpec (MaskSpec), MAE-AST (mae-ast) and Audio-MAE (audiomae). Of the masked auto-encoder approaches, MW-MAE (mwmae) achieves state-of-the-art performance on the HEAR benchmark by using multi-window local-global attention in the decoder. Recently, SSAM (audiomamba) utilized a Mamba (mamba) architecture in their encoder and achieved similar performance as MW-MAE. In contrast to the masked auto-encoders, BEATS (beats) utilizesmasking-based approach based on latent embeddings extracted by an acoustic tokenizer. Finally, successful self-supervised approaches that do not rely on masking at all include contrastive learning frameworks such as COLA (COLA).
Another category of self-supervised audio representation models focuses on speech representations specifically, making use of generative, predictive or contrastive learning (mohamed2022self). These speech models are typically trained on datasets such as Librispeech (librispeech) or LibriLight (librilight) and include state-of-the-art models such as Wav2Vec2 (wav2vec2), HuBERT (hubert) and WavLM (wavlm). However, while these models excel at speech-based tasks, they do not necessarily generalize well to non-speech sounds and non-speech tasks (HEAR). Crucially, none of the existing self-supervised approaches for audio or speech representation learning optimize for performance in real-world sound scenes that are spatial, reverberant, and noisy.
@@ 3.1 Simulating real-world acoustic scenes
Pipeline overview: A room impulse response (RIR) captures room specific acoustic properties such as reverberation. We utilized high-resolution, detailed 3D meshes of houses with various architectural characteristics from Matterport3D (Matterport3D) in order to simulate RIRs for many different rooms in each house with the Monte Carlo ray tracing RIR simulator provided by SoundSpaces 2.0 chen22soundspaces2. SoundSpaces 2.0 combines the simulated RIRs with a head-related transfer function (HRTF) (CIPIC) to generate a binaural RIR (BRIR) or with an ambisonics microphone configuration to generate an ambisonics RIR (ARIR). BRIRs capture both room acoustic properties and human spatial hearing characteristics introduced by the shape of the ears, head and torso, while ARIRs capture room acoustic properties as well as the spatial cues encoded in first-order Ambisonics.
@@ 3.2 GRAM framework
The GRAM learns spatial audio representation by reconstructing multi-channel masked spectrogram patches. First, a patch extractor consisting of a single convolutional layer with 2D convolutional filters divides each multi-channel spectrogram into $n$ non-overlapping patches $P_{1},\ldots,P_{n}$ with $P_{i}\in\mathbb{R}^{C\times T\times F}$, and embeds each patch into a linear patch embedding $E_{i}\in\mathbb{R}^{768}$ (Figure 1). Non-masked patch embeddings are input to the encoder, for which we selected the 12-layer ViT-Base (ViT-B) Transformer (dosovitskiy2020vit) similar to audiomae; mwmae. The encoder outputs patch representations $O_{i}\in\mathbb{R}^{768}$ for $i=1,\ldots,n$, where $n$ is the number of unmasked patches. Finally, a Transformer decoder with local-global attention (mwmae) followed by a linear head takes all patch representations $O_{1},\ldots,O_{n}$ as well as all masked patches $M_{1},\ldots,M_{n}$ to reconstruct the multi-channel spectrogram from last layer embeddings.
@@ 3.3 Pre-training
Online mixing of naturalistic sound scenes: The 85,000 naturalistic scenes were split into a train set of 70,000 scenes (corresponding to 70 Matterport3D houses), and a test set of 15,000 scenes (15 Matterport3D houses) for down-stream evaluation (see Section 3.4). We used the 70,000 naturalistic scenes in the train set to generate naturalistic scenes for all audio clips in the unbalanced training set of AudioSet (10-second sound tracks of 1.74 million YouTube videos (audioset)). Specifically, during training we randomly paired an AudioSet clip with a noise sound clip from the WHAMR! background noise database (Maciejewski2020WHAMR). WHAMR! noise clips longer than 10 s were trimmed to 10 s duration and a linear fade-in/fade-out of 200 ms was applied to every noise clip prior to mixing of the sound scene.
@@
Patch extraction and masking: For pre-training, we divided the binaural spectrogram into $P_{i}\in\mathbb{R}^{2\times 8\times 16}$, and ambisonics spectrograms into $P_{i}\in\mathbb{R}^{7\times 8\times 16}$ patches. We used an adapted version of the mask-based framework of MW-MAE (mwmae), randomly selecting a subset of $n$ patches $M_{1},\ldots,M_{n}$ for $i=1,\ldots,n$ for masking (masking ratio = 0.8) and replacing their embedding with a learnable mask token. Finally, we added fixed sinusoidal positional embeddings to all embedded patches.
Decoder with local-global attention: The decoder takes as input both the unmasked patches $O_{1},\ldots,O_{n}$ with $O_{i}\in\mathbb{R}^{768}$, and the masked patches $M_{1},\ldots,M_{n}$ with $M_{i}\in\mathbb{R}^{768}$ as well as fixed sinusoidal positional embeddings for each patch (Figure 1). To implement local-global attention (mwmae), we selected window sizes of [2, 5, 10, 25, 50, 100, 0, 0]. Here, 0 signifies global plain attention.
Pre-training specification: We trained all GRAMs for 500 K steps on an H100 92 GB GPU machine with 16 CPU cores. We used the AdamW optimizer (loshchilov2017decoupled) with weight decay rate of 0.01, gradient clipping, and a cosine learning rate scheduler with 10 K steps warm-up. The initial learning rate was set to 0.0002, and decayed to 0. We optimize the mean squared error (MSE) loss function between the predicted masked patches and their corresponding input spectrogram patches.
@@ 3.4 Experiments
Model comparison: We compare the performance and efficiency of GRAM-Binaural, GRAM-Ambisonics and GRAM-Clean (trained on regular AudioSet) on downstream tasks with state-of-the-art self-supervised audio representation models with a similar number of parameters as GRAM (90 M): MAE-16x16 (audiomae), SSAST-patch (gong2022ssast), BEATs-iter3 (beats), MW-MAE-B-200-4x16 (mwmae), SSAM (audiomamba); self-supervised speech representation models Wav2Vec 2.0 Base (wav2vec2), HuBERT Base (hubert), WavLM Base (wavlm).
@@
Downstream tasks: We evaluate GRAM and other state-of-the-art models on the HEAR benchmark task suite, which presents a wide range of tasks to evaluate the downstream performance of audio representation models (HEAR). To avoid redundancy we selected the same subset of HEAR tasks as previously used in (mwmae). To enable in-depth evaluation of audio scene analysis capabilities, we added the time-stamp based sound event detection task DCASE-2016 Task 2 (Mesaros2018_TASLP) from the HEAR benchmark suite.
@@
We additionally evaluated performance on simulated real-worldsound scenes using Nat-HEAR, which provides a naturalistic version of all selected datasets in the HEAR benchmark suite in two audio formats: a two-channel, binaural format and a four-channel, first-order Ambisonics format. Weincluded sound localization tasks for two different domains which we generated using HEAR benchmark datasets: A speech localization task based on SC-5, and an environmental sound localization task based on ESC-50. The localization tasks are modeled as a multi-output regression task in which model outputs represent the estimated 3D Cartesian coordinates $[x,y,z]$ on the unit sphere (adavanne2018sound). Finally, to assess the transferability of GRAM to real-world sound scenes, we evaluate also on the sound event detection and localization tasks in TUT Sound Events 2018 REAL (tut).
@@
Downstream evaluation: To evaluate the single-channel SOTA audio representation models on Nat-HEAR, we utilized the omnidirectional channel $W$ of the first-order Ambisonics (ambisonics) version of Nat-HEAR as model input. The outputs of MAE, SSAST and BEATs were not suitable for the time-stamp based DCASE-2016 sound event detection task. Hence, these models were not evaluated on the DCASE-2016 task. Further, we included Spatial-AST (spatialast) as this is the only model trained on spatial sound scenes and therefore the sole model evaluated on the Nat-HEAR localization tasks besides GRAM-Binaural and GRAM-Ambisonics. To evaluate two-channel models such as GRAM-Binaural on HEAR, we duplicated the single-channel spectrograms of the original HEAR to generate model compatible input. Following the HEAR protocol (HEAR) for downstream task evaluation, we extracted embeddings from the frozen pretrained models and trained a shallow downstream classifier on these embeddings to assess how well the learned representations generalize to a broad range of tasks. The embedding extraction for GRAM is described in Appendix D.
@@
Figure 2: Downstream model performance on naturalistic sound scenes. (A) Nat-HEAR and HEAR downstream performance as a function of training data quantity. (B) Box plots of the difference in performance on HEAR and Nat-HEAR, excluding the DCASE-2016 task. Box limits reflect the first and third quartile, center line the median (see also Table 5).
@@
Quantifying overall performance: We calculate for each model $m$, the score $s(m)$ to give an impression of the overall performance, similar to (superb). This score reflects a model’s improvement with respect to the maximum improvement over the baseline obtained by the current state-of-the-art model, averaged across all tasks included in the benchmark. This metric effectively ranks the improvement of models over the baseline as a function of the current maximum improvement (see Appendix D). We use the HEAR-Naive baseline based on mel-spectrograms (HEAR).
@@
Generalization to recorded real-world scenes: To evaluate GRAM-Ambisonics on the sound event detection and sound localization takss of TUT Sound Events 2018 REAL (tut), resampled waveforms to 32 kHz and extracted segments with their respective localization labels. We modeled the localization task as polar coordinate regression, predicting [azimuth, elevation], and used the same fine-tuning protocol as for HEAR tasks.